OpenStudy (anonymous):

what is the area, in sq millimeters, of a right triangle with a hypotenuse of 68 millimeters and a leg of 60 millimeters? a. 2,040 b. 1,920 c. 1,088 d. 960

OpenStudy (smarty):

B 32*60=1920

OpenStudy (anonymous):

See, hypotenuse is given and one of its side given can be thought as Either base or Height.. Let We suppose that Height is given as 60 mm.. Now, \[Base = \sqrt{(Hypotenuse^2 - Height^2)}\] So, base will be :\[B = \sqrt{68^2 - 60^2} = \sqrt{(4624 - 3600)} = \sqrt{1024} = 32\] So Base = 32 mm Now use:\[Area =(Base)*(Height)/2 = 32*60/2 = 32*30 = 960 mm\] is the answer.. So, your option is D...